FIND THE INTEGRATION OF  elog(1+tan^2x) xdx ?
 
 AND ILLUSTRATE HOW IT WILL COME ?
 

Grade:12th Pass

1 Answers

jagdish singh singh
8 years ago
\hspace{-0.5cm}$Let $\bf{I = \int e^{\ln(1+\tan^2 x)}dx = \int (1+\tan^2 x)dx = \int \sec^2 xdx = \tan x+\mathcal{C}.}$\\\\ Above we have use the formula $\bullet\; \bf{e^{\ln (x)} = x.}$ \\\\